What is the length and width of a rectangle with 80 square meters and the length of the rectangle is 2 meters more then its width?
Answer by checkley75(3666)

L=W+2
LW=80
(W+2)W=80
W^2+2W-80=0
(W+10)(W-8)=0
W-8=0
W=8 ANS. FOR THE WIDTH.
L=8+2=10 ANS. FOR THE LENGTH.
PROOF:
8*10=80
80=80